# Recursive Training Loops in LLMs: How training data properties modulate distribution shift in generated data?

### v5

@@ 1 Introduction
Large Language Models (LLMs) are increasingly contributing to the creation of internet content, being used for journalism (Brigham et al., 2024), coding Jiang et al. (2024) and generating content on social media Ferrara et al. (2016). The increasing amount of synthetic, LLM-generated data on the internet introduces a precarious feedback loop: LLMs trained on datasets containing synthetic data will themselves generate data that will be used to train future models. Shumailov et al. (2024) demonstrated that this process, known as recursive training, can have detrimental effects, causing models to progressively lose information about the true underlying distributions they are intended to approximate. This results in a gradual change in generated distributions, often accompanied by a reduction in variance. While this has sometimes been described as model collapse, we refer to this mismatch between the true and the generated distribution as distribution shift, and limit the term collapse to refer to detrimental distribution shifts, such as losses in quality or diversity. Such detrimental effects Guo et al. (2024) as well as bias amplification Wang et al. (2024a) have been reported in previous work. Their potential societal consequences make it imperative to better understand the dynamics of recursive training.
@@
Given the early stage of research in this field, existing studies often ignore this diversity, relying on simplifying assumptions and focusing on abstracted settings. To the best of our knowledge, how different data properties influence distribution shifts remains largely unexplored, aside from studies examining the ratio of human to synthetic data (Bertrand et al., 2023; Bohacek and Farid, 2023; Kazdan et al., 2024; Martínez et al., 2023b; Zhang et al., 2024b). Filing this gap is therefore crucial to draw a more nuanced and detailed picture of the consequences of recursive fine-tuning.
In this paper, we adopt the iterative chain paradigm used in previous studies (Shumailov et al., 2024; Gerstgrasser et al., 2024a). Our experimental setup is shown in Figure 1. The process begins with fine-tuning a base LLM on a selection of human data (e.g. Reddit posts). This model generates data, which are added to the pool of Accumulated data together with newly sampled human data. In each subsequent generation, a new base model is fine-tuned on data sampled from the Accumulated data pool. This model, in turn, generates new data, which are again added to the pool along with fresh human data. This pipeline allows us to study how generated data evolve across successive generations, with a particular focus on the distribution shift from the first to the final generation.
@@ 2 Related Work
A rapidly growing body of literature has focused on the consequences of recursively training generative models on synthetic data Schaeffer et al. (2025). The phrase “model collapse”, coined in Shumailov et al. (2024), refers to the progressive degradation of models induced by this feedback loop. This phenomenon has been studied both theoretically (Dohmatob et al., 2024a, b; Bertrand et al., 2023; Alemohammad et al., 2023) and empirically, on both generative image models (Martínez et al., 2023b, a; Bohacek and Farid, 2023; Hataya et al., 2022; Alemohammad et al., 2023) and language models Zhang et al. (2024b); Guo et al. (2023); Kazdan et al. (2024); Briesch et al. (2023); Gerstgrasser et al. (2024b). Theoretical results have provided valuable insights, for instance showing how it is characterized by the disappearance of distribution tails Dohmatob et al. (2024a, b); Shumailov et al. (2024). Empirical studies have allowed to establish several properties of this phenomenon, such as the role of synthetic-to-real-data ratio (Briesch et al., 2023; Hataya et al., 2022) or strategies for mitigating collapse (Gerstgrasser et al., 2024b; Kazdan et al., 2024; Zhang et al., 2024b). Recently, Wang et al. (2024b) showed that recursive LLM fine-tuning can lead to bias amplification. These works do not systematically evaluate how the properties of the human dataset used in their experiments affect their conclusions. Here, we extend this literature by investigating how those properties impact the outcome of recursive training.
@@
The motivation for this research area stems from the observation that human-made technologies have transitioned from passive mediators of cultural evolution (e.g., the printing press) to active generators of cultural content. This shift has been described as the emergence of machine culture - culture mediated or generated by machines Brinkmann et al. (2023). Understanding the dynamics that shape the evolution of machine-generated content over time is therefore crucial. This has led researchers to study cultural dynamics in populations of reinforcement learning agents (Cook et al., 2024; Schmitt et al., 2018; Team et al., 2021; Prystawski et al., 2023; Nisioti et al., 2022) and of LLMs Perez et al. (2024a, b); Nisioti et al. (2024); Vallinder and Hughes (2024); Burton et al. (2024). Our work extends this literature by examining the factors that modulate the evolution of LLM-generated content.
@@ 3.1 The iterative chain paradigm
We use the iterative chain paradigm inspired by Shumailov et al. (2024); Gerstgrasser et al. (2024a). Our experimental design is shown in Figure 1. First, a base LLM is fine-tuned on 8000 samples from a human dataset (e.g. Wikipedia articles or Reddit posts). This model generates 4000r4000*r posts, where $r$ is the synthetic-data ratio. Those posts are added to the pool of accumulated data together with 4000(1r)4000*(1-r) newly sampled human posts. In all subsequent generations, a new base model is fine-tuned on $4000$ posts sampled from the accumulated data pool, and 4000r4000*r posts generated by this model are added to the accumulated data pool together with 4000(1r)4000*(1-r) newly sampled human data. In each generation, a new base model is sampled from four possible options: LLaMa-3.2-1B (Dubey et al., 2024) (LLAMA 3.1 Community), Qwen2.5-1.5B (Team, 2024b) (Apache), SmolLM-1.7B (Allal et al., 2024) (Apache), Falcon3-1B-Base (Team, 2024a)), and fine-tuned using LoRA (Hu et al., 2021) (see appendix B.5 for details). Five seeds were used in all experiments. This pipeline enables us to study the evolution of generated data over generations, and most notably, the difference between data generated in the first and last generations.
@@ 3.3 Metrics
The Semantic Diversity of a set of texts is measured as the pairwise cosine diversity in the stella_en_1.5B_v5 model (Zhang et al., 2024a), as in previous studies (Guo et al., 2024). Quality and Political lean are estimated by using LLaMa-3.3-70B-Instruct (Dubey et al., 2024) in the LLM-as-a-Judge setup. The prompt is inspired by Wang et al. (2023) and Chen et al. (2023), and adapted to our task of evaluating the quality of short texts (see appendix B.3 or details and validation of that adaptation). In sections 4.3 and 4.4, we additionally rely on the following metrics. Lexical Diversity is estimated as SelfBLEU (Zhu et al., 2018), computed as the average BLEU score Papineni et al. (2002) of each text using all other texts as references, following prior work Guo et al. (2024). Gaussianity is measured by fitting a 2D UMAP projection on embeddings from the stella_en_1.5B_v5 model, and computing the AIC (Akaike, 1974) of a 2D Gaussian distribution fit to this space. Positivity is assessed using the SentimentIntensityAnalyzer tool from NLTK (Hardeniya et al., 2016), which assigns a sentiment score to each text ranging from 1.0-1.0 (highly negative) to $1.0$ (highly positive).
@@ 4.1 Does synthetic data ratio impact distribution shift dynamics?
In this experiment, we aim to reproduce the effect of the synthetic data ratio (Briesch et al., 2023) on the shifts toward lower quality and lower diversity induced by recursive fine-tuning. This ratio corresponds to the proportion between the AI-generated data and the freshly sampled human data that are added to the Accumulated data pool at each generation.
@@ B.2 Dataset details
($C$-BY-4.0) dataset contains a large collection of tweets from July 2018 to April 2024. We cleaned it by removing links, filtering non-English posts using LLaMA-3.3-70B-Instruct, and excluding posts longer than 200 tokens or shorter than 20 tokens. Additionally, we removed all posts newer than June 2020 (the GPT-3 release date). The final cleaned dataset consists of 2 million posts. The senator_tweets  [^3] https://huggingface.co/datasets/m-newhauser/senator-tweets dataset contains all tweets made by United States senators during the first year of the Biden Administration (2021). We cleaned it by removing links and posts shorter than 10 tokens. The final cleaned dataset consists of 94878 posts. The wikipedia (Foundation, )  [^4] https://huggingface.co/datasets/wikimedia/wikipedia (CC-BY-SA-3.0) dataset was created by compiling and cleaning articles from Wikipedia dumps  [^5] https://dumps.wikimedia.org/ in November 2023. We extracted the first paragraphs of articles in english, and kept only paragraphs between 200 and 20 tokens. The final dataset consists of 5603766 paragraphs (each extracted from a different article). The reddit_submissions  [^6] https://huggingface.co/datasets/HuggingFaceGECLM
@@
(arXiv.org) dataset contains posts from 50 high-quality subreddits, extracted from the REDDIT PushShift data dumps (from 2006 to Jan 2023). We pre-processed this dataset by merging post titles with bodies, sampling 25000 posts from each Subreddit, removing those that have [deleted] or [removed] tags, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1243794 posts. The webis_reddit Völske et al. (2017) (CC-BY) dataset contains preprocessed posts from the Reddit dataset (Webis-TLDR-17). We pre-processed this dataset by merging titles with bodies, removing "tldr" tags, removing posts that are marked as "nsfw" or "+18", removing duplicates, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1458003 posts.
@@ B.3 LLM-as-a-judge validation
We measured text quality using LLM-as-a-judge method whose performance been empirically confirmed in previous studies Chen et al. (2023). We use LLama-3.3-70B-Instruct to annotate texts on a scale of 0 to 100 using the following prompt:
@@
To confirm the correctness of our judge, we compare it to the judge from Chen et al. (2023), which was shown to outperform many other quality metrics. We use the same experimental setup as in Chen et al. (2023) and compare to numbers extracted from Table 10 in Chen et al. (2023). Table 2 shows the correlations for the two judges with human annotations of the OpenMEVA-ROC dataset Guan et al. (2021) of generated stories. We can see that our judge matches that of Chen et al. (2023) justifying its use in our experiments.
@@
Chen et al. (2023) 53.5 50.8
Table 2: Comparison of our judge to that of Chen et al. (2023) using human annotations on OpenMEVA-ROC Guan et al. (2021). Our judge matches their performance justifying its use in this paper.
@@ B.3.1 Data properties selection
  • Semantic Diversity: We compute several semantic diversity metrics as pairwise cosine diversity in the embeddings of the stella_en_1.5B_v5 model (Zhang et al., 2024a). Cosine diversity computes the pairwise diversity between all data points, and k-nn Cosine diversity computes the pairwise diversity for the nearest n neighbors (we use this metric with n=50 and n=1000).
  • Lexial diversity: Self-BLEU (Zhu et al., 2018) is a metric that computes the average BLEU(Papineni et al., 2002) score for each text, with all other text taken as references.
@@
  • Type Token Ratio (TTR) (Johnson, 1944): calculates the number of unique words (types) divided by the number of total word in the first 200 characters of each text.
@@
  • Positivity: uses the SentimentIntensityAnalyzer tool from NLTK Hardeniya et al. (2016))(Apache) to assign a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
  • Toxicity: quantifies the presence of rude, disrespectful, or unreasonable language, using a probability score that ranges from 0.0 (benign and non-toxic) to 1.0 (highly likely to be toxic), as estimated by the classifier introduced in Hanu and Unitary team (2020).
  • KL-Entropy: fits a 2D UMAP on the stella_en_1.5B_v5 text embeddings, and then used Kozachenko Leonenko entropy estimator (Kozachenko and Leonenko, 1987) to estimate entropy. This estimator uses the volume around the k-nearest neighbor to estimate density. We create 2 kl_entropy metics, one with k=50 and one with k=1000.
  • Gaussianity: uses the same UMAP representations as KL-entropy. In this space, it fits a 2D Gaussian distribution. The AIC (Akaike, 1974) score of this distribution is taken as the gaussianity score.
@@ B.4 Clustering
To obtain data to be used for regression analysis, we create a number of subsets from each of the datasets. This is done by the following procedure. A 2D UMAP is fit on the 90k text embedded with stella_en_1.5B_v5 model (Zhang et al., 2024a). A series of clustering methods (dbscan, hdbscan, gmm, k-means) with different hyperparameters are done separately, each annotating the 90k samples. Then the rest of the dataset is annotated by k-nn, with k=1. This is done in two ways, with and without excluding the noise cluster for the k-nn classifier. This gives us a total of 120 different clusterings of the dataset. Then for each clustering 10 clusters are taken, if there are not enough clusters over the size of 60k. The remaining clusters are constructed by merging smaller clusters, either by uniformly sampling which cluster to merge or by iteratively merging the cluster that is the furthest away from the currently merged cluster. As this results in a large number of clusters (e.g. 1088 for webis_reddit ), we then obtain the final cluster set by subsampling 200 clusters.
@@ B.5 Fine-tuning procedure
We use the Unsloth library (Daniel Han and team, 2023)(Apache license) to train model suing LoRA (Hu et al., 2021). The hyperparameters used are the default ones given by Unsloth, i.e. rank = 16, alpha = 16 batch size = 16, leaning rate = 2e-4, we use a linear schedules with 5 steps of warm-up. For generation with use a temperature of 1.5 with min_p (Nguyen et al., 2024) of 0.2.

### v6

@@ 1 Introduction
Large Language Models (LLMs) are increasingly contributing to the creation of internet content, being used for journalism (llm_news), coding llm_code_survey and generating content on social media the_rise_of_social_bots. The increasing amount of synthetic, LLM-generated data on the internet introduces a precarious feedback loop: LLMs trained on datasets containing synthetic data will themselves generate data that will be used to train future models. model_collapse demonstrated that this process, known as recursive training, can have detrimental effects, causing models to progressively lose information about the true underlying distributions they are intended to approximate. This results in a gradual change in generated distributions, often accompanied by a reduction in variance. While this has sometimes been described as model collapse, we refer to this mismatch between the true and the generated distribution as distribution shift, and limit the term collapse to refer to detrimental distribution shifts, such as losses in quality or diversity. Such detrimental effects curious_decline_div as well as bias amplification bias_amplification have been reported in previous work. Their potential societal consequences make it imperative to better understand the dynamics of recursive training.
@@
Given the early stage of research in this field, existing studies often ignore this diversity, relying on simplifying assumptions and focusing on abstracted settings. To the best of our knowledge, how different data properties influence distribution shifts remains largely unexplored, aside from studies examining the ratio of human to synthetic data (Bertrand2023; Bohacek2023; Kazdan2024; Martnez2023; Zhang2024). Filing this gap is therefore crucial to draw a more nuanced and detailed picture of the consequences of recursive fine-tuning.
In this paper, we adopt the iterative chain paradigm used in previous studies (model_collapse; accumulation). Our experimental setup is shown in Figure 1. The process begins with fine-tuning a base LLM on a selection of human data (e.g. Reddit posts). This model generates data, which are added to the pool of Accumulated data together with newly sampled human data. In each subsequent generation, a new base model is fine-tuned on data sampled from the Accumulated data pool. This model, in turn, generates new data, which are again added to the pool along with fresh human data. This pipeline allows us to study how generated data evolve across successive generations, with a particular focus on the distribution shift from the first to the final generation.
@@ 2 Related Work
A rapidly growing body of literature has focused on the consequences of recursively training generative models on synthetic data schaeffer2025position. The phrase “model collapse”, coined in model_collapse, refers to the progressive degradation of models induced by this feedback loop. This phenomenon has been studied both theoretically (Dohmatob2024_demist; dohmatob2024tale; Bertrand2023; Alemohammad2023) and empirically, on both generative image models (Martnez2023; Martnez2023b; Bohacek2023; Hataya2022; Alemohammad2023) and language models Zhang2024; Guo2023; Kazdan2024; Briesch2023; Gerstgrasser2024. Theoretical results have provided valuable insights, for instance showing how it is characterized by the disappearance of distribution tails Dohmatob2024_demist; dohmatob2024tale; model_collapse. Empirical studies have allowed to establish several properties of this phenomenon, such as the role of synthetic-to-real-data ratio (Briesch2023; Hataya2022) or strategies for mitigating collapse (Gerstgrasser2024; Kazdan2024; Zhang2024). Recently, Wang2024 showed that recursive LLM fine-tuning can lead to bias amplification. These works do not systematically evaluate how the properties of the human dataset used in their experiments affect their conclusions. Here, we extend this literature by investigating how those properties impact the outcome of recursive training.
@@
The motivation for this research area stems from the observation that human-made technologies have transitioned from passive mediators of cultural evolution (e.g., the printing press) to active generators of cultural content. This shift has been described as the emergence of machine culture - culture mediated or generated by machines machine_culture. Understanding the dynamics that shape the evolution of machine-generated content over time is therefore crucial. This has led researchers to study cultural dynamics in populations of reinforcement learning agents (cook_artificial_2024; schmitt_kickstarting_2018; open_ended_learning_team_open-ended_2021; prystawski_cultural_2023; nisioti_social_2022) and of LLMs perez2024llms; perez2024cultural; nisioti2024collective; vallinder2024cultural; burton2024large. Our work extends this literature by examining the factors that modulate the evolution of LLM-generated content.
@@ 3.1 The iterative chain paradigm
We use the iterative chain paradigm inspired by model_collapse; accumulation. Our experimental design is shown in Figure 1. First, a base LLM is fine-tuned on 8000 samples from a human dataset (e.g. Wikipedia articles or Reddit posts). This model generates 4000r4000*r posts, where $r$ is the synthetic-data ratio. Those posts are added to the pool of accumulated data together with 4000(1r)4000*(1-r) newly sampled human posts. In all subsequent generations, a new base model is fine-tuned on $4000$ posts sampled from the accumulated data pool, and 4000r4000*r posts generated by this model are added to the accumulated data pool together with 4000(1r)4000*(1-r) newly sampled human data. In each generation, a new base model is sampled from four possible options: LLaMa-3.2-1B (llama_3) (LLAMA 3.1 Community), Qwen2.5-1.5B (qwen2.5) (Apache), SmolLM-1.7B (smollm) (Apache), Falcon3-1B-Base (falcon3)), and fine-tuned using LoRA (lora) (see appendix B.5 for details). Five seeds were used in all experiments. This pipeline enables us to study the evolution of generated data over generations, and most notably, the difference between data generated in the first and last generations.
@@ 3.3 Metrics
The Semantic Diversity of a set of texts is measured as the pairwise cosine diversity in the stella_en_1.5B_v5 model (stella), as in previous studies (curious_decline_div). Quality and Political lean are estimated by using LLaMa-3.3-70B-Instruct (llama_3) in the LLM-as-a-Judge setup. The prompt is inspired by is_chatgpt_a_good_nlg_evaluator and llm_for_reference_free_quality_eval, and adapted to our task of evaluating the quality of short texts (see appendix B.3 or details and validation of that adaptation). In sections 4.3 and 4.4, we additionally rely on the following metrics. Lexical Diversity is estimated as SelfBLEU (selfbleu), computed as the average BLEU score bleu of each text using all other texts as references, following prior work curious_decline_div. Gaussianity is measured by fitting a 2D UMAP projection on embeddings from the stella_en_1.5B_v5 model, and computing the AIC (aic) of a 2D Gaussian distribution fit to this space. Positivity is assessed using the SentimentIntensityAnalyzer tool from NLTK (hardeniya2016natural), which assigns a sentiment score to each text ranging from 1.0-1.0 (highly negative) to $1.0$ (highly positive).
@@ 4.1 Does synthetic data ratio impact distribution shift dynamics?
In this experiment, we aim to reproduce the effect of the synthetic data ratio (Briesch2023) on the shifts toward lower quality and lower diversity induced by recursive fine-tuning. This ratio corresponds to the proportion between the AI-generated data and the freshly sampled human data that are added to the Accumulated data pool at each generation.
@@ B.2 Dataset details
($C$-BY-4.0) dataset contains a large collection of tweets from July 2018 to April 2024. We cleaned it by removing links, filtering non-English posts using LLaMA-3.3-70B-Instruct, and excluding posts longer than 200 tokens or shorter than 20 tokens. Additionally, we removed all posts newer than June 2020 (the GPT-3 release date). The final cleaned dataset consists of 2 million posts. The senator_tweets  [^3] https://huggingface.co/datasets/m-newhauser/senator-tweets dataset contains all tweets made by United States senators during the first year of the Biden Administration (2021). We cleaned it by removing links and posts shorter than 10 tokens. The final cleaned dataset consists of 94878 posts. The wikipedia (wikipedia_dataset)  [^4] https://huggingface.co/datasets/wikimedia/wikipedia (CC-BY-SA-3.0) dataset was created by compiling and cleaning articles from Wikipedia dumps  [^5] https://dumps.wikimedia.org/ in November 2023. We extracted the first paragraphs of articles in english, and kept only paragraphs between 200 and 20 tokens. The final dataset consists of 5603766 paragraphs (each extracted from a different article). The reddit_submissions  [^6] https://huggingface.co/datasets/HuggingFaceGECLM
@@
(arXiv.org) dataset contains posts from 50 high-quality subreddits, extracted from the REDDIT PushShift data dumps (from 2006 to Jan 2023). We pre-processed this dataset by merging post titles with bodies, sampling 25000 posts from each Subreddit, removing those that have [deleted] or [removed] tags, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1243794 posts. The webis_reddit webis (CC-BY) dataset contains preprocessed posts from the Reddit dataset (Webis-TLDR-17). We pre-processed this dataset by merging titles with bodies, removing "tldr" tags, removing posts that are marked as "nsfw" or "+18", removing duplicates, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1458003 posts.
@@ B.3 LLM-as-a-judge validation
We measured text quality using LLM-as-a-judge method whose performance been empirically confirmed in previous studies llm_for_reference_free_quality_eval. We use LLama-3.3-70B-Instruct to annotate texts on a scale of 0 to 100 using the following prompt:
@@
To confirm the correctness of our judge, we compare it to the judge from llm_for_reference_free_quality_eval, which was shown to outperform many other quality metrics. We use the same experimental setup as in llm_for_reference_free_quality_eval and compare to numbers extracted from Table 10 in llm_for_reference_free_quality_eval. Table 2 shows the correlations for the two judges with human annotations of the OpenMEVA-ROC dataset guan-etal-2021-openmeva of generated stories. We can see that our judge matches that of llm_for_reference_free_quality_eval justifying its use in our experiments.
@@
llm_for_reference_free_quality_eval 53.5 50.8
Table 2: Comparison of our judge to that of llm_for_reference_free_quality_eval using human annotations on OpenMEVA-ROC guan-etal-2021-openmeva. Our judge matches their performance justifying its use in this paper.
@@ B.3.1 Data properties selection
  • Semantic Diversity: We compute several semantic diversity metrics as pairwise cosine diversity in the embeddings of the stella_en_1.5B_v5 model (stella). Cosine diversity computes the pairwise diversity between all data points, and k-nn Cosine diversity computes the pairwise diversity for the nearest n neighbors (we use this metric with n=50 and n=1000).
  • Lexial diversity: Self-BLEU (selfbleu) is a metric that computes the average BLEU(bleu) score for each text, with all other text taken as references.
@@
  • Type Token Ratio (TTR) (ttr): calculates the number of unique words (types) divided by the number of total word in the first 200 characters of each text.
@@
  • Positivity: uses the SentimentIntensityAnalyzer tool from NLTK hardeniya2016natural)(Apache) to assign a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
  • Toxicity: quantifies the presence of rude, disrespectful, or unreasonable language, using a probability score that ranges from 0.0 (benign and non-toxic) to 1.0 (highly likely to be toxic), as estimated by the classifier introduced in Detoxify.
  • KL-Entropy: fits a 2D UMAP on the stella_en_1.5B_v5 text embeddings, and then used Kozachenko Leonenko entropy estimator (kl_ent) to estimate entropy. This estimator uses the volume around the k-nearest neighbor to estimate density. We create 2 kl_entropy metics, one with k=50 and one with k=1000.
  • Gaussianity: uses the same UMAP representations as KL-entropy. In this space, it fits a 2D Gaussian distribution. The AIC (aic) score of this distribution is taken as the gaussianity score.
@@ B.4 Clustering
To obtain data to be used for regression analysis, we create a number of subsets from each of the datasets. This is done by the following procedure. A 2D UMAP is fit on the 90k text embedded with stella_en_1.5B_v5 model (stella). A series of clustering methods (dbscan, hdbscan, gmm, k-means) with different hyperparameters are done separately, each annotating the 90k samples. Then the rest of the dataset is annotated by k-nn, with k=1. This is done in two ways, with and without excluding the noise cluster for the k-nn classifier. This gives us a total of 120 different clusterings of the dataset. Then for each clustering 10 clusters are taken, if there are not enough clusters over the size of 60k. The remaining clusters are constructed by merging smaller clusters, either by uniformly sampling which cluster to merge or by iteratively merging the cluster that is the furthest away from the currently merged cluster. As this results in a large number of clusters (e.g. 1088 for webis_reddit ), we then obtain the final cluster set by subsampling 200 clusters.
@@ B.5 Fine-tuning procedure
We use the Unsloth library (unsloth)(Apache license) to train model suing LoRA (lora). The hyperparameters used are the default ones given by Unsloth, i.e. rank = 16, alpha = 16 batch size = 16, leaning rate = 2e-4, we use a linear schedules with 5 steps of warm-up. For generation with use a temperature of 1.5 with min_p (minp) of 0.2.